# Does Privacy Always Harm Fairness? Data-Dependent Trade-offs via Chernoff Information Neural Estimation

### v1

@@ B.1.1 Tightness of Bound
where $n$ is the number of samples (Cover, 1999). For instance, for Gaussian distributions, we can explicitly write:
@@ B.2 Connection to Fairness Metrics
Lemma B.2 (Lemma 2 from Dutta et al. (2020)).
@@
Definition B.1 (Chernoff Exponents—Definition 1 from Dutta et al. (2020)).
@@
Lemma B.3 (Chernoff Bound - Lemma 1 from Dutta et al. (2020)).
@@
Definition B.2 (Definition 2 from Dutta et al. (2020)).
@@
Definition B.3 (Section E.1 (Dutta et al., 2020)).
@@ Appendix C Isotropic Gaussian Proofs
Lemma C.1 (Restatement of Lemma 3).
@@ Appendix D Supplemental Gaussian Figures
D.1 Fairness-Accuracy vs Log Fairness-Accuracy
@@
D.2 Case 2: More detail
@@ E.2 Proofs
Computing the infimum over the open interval (0,1) and closed interval [0,1] yield the same result following the strict convexity of $f$. We point to Remark 2 from Nielsen (2022) for a discussion of the boundary conditions. In fact, in some references, Chernoff Information is presented using this closed interval (Cover, 1999). This provides a compactness property that can be utilized in the following lemmas. ∎
Lemma E.2.
@@
This follows from Theorem 1 from Newey (1991) as well as pointwise convergence in Lemma E.2 and stochastic equicontinuity in Lemma E.3. ∎
@@
This result follows from Theorem E.1. Continuity and the uniqueness of the maximizer are given by Nielsen (2022). Compactness is given by Lemma E.1. Uniform Convergence in probability is given by Lemma E.4. Hence,
@@ F.1.1 Time Complexity
where $n$ is the number of samples used to compute Chenroff Information, $T_{\text{DRE}}$ is time spent on density ratio estimation (DRE) in Line 1, $C_{\text{int}}$ is the time required for one integral operations given in Line 2, $C_{MC}$ is the time required for simple addition/floating point operations required in Line 3, and $T_{\text{cvx}}$ is the time required to solve the convex optimization in Line 4, which is a constant time operation. Hence, for other than $T_{\text{DRE}}$, the rest of the operation is done in $O(n)$ and training a neural network for DRE is often the most time-consuming part, which took approximately 30 minutes on a single NVIDIA L40S GPU.
F.2 Gaussian
@@ F.3 Datasets
We replicate the previous experiments using 2D embeddings of MNIST data which were obtained via Principal Component Analysis. For $P_{0}$, we select the digit 0, $P_{1}$, we select the digit $2$, $Q_{0}$, we select the digit $3$ and $Q_{1}$ we select the digit 7.

### v2

@@ B.1.1 Tightness of Bound
where $n$ is the number of samples (Cover and Thomas, 1999). For instance, for Gaussian distributions, we can explicitly write:
@@ B.2 Connection to Fairness Metrics
Lemma 7 (Lemma 2 from Dutta et al. (2020)).
@@
Definition 9 (Chernoff Exponents—Definition 1 from Dutta et al. (2020)).
@@
Lemma 8 (Chernoff Bound - Lemma 1 from Dutta et al. (2020)).
@@
Definition 10 (Definition 2 from Dutta et al. (2020)).
@@
Definition 11 (Section E.1 (Dutta et al., 2020)).
@@
B.3 Multi-group Fairness
While we focus on the binary setting, we can generalize our framework to multi-group settings. In these settings, Chernoff Information can play the role of a group statistic (analogous to group accuracy or group false-positive rate), which can be compared across all groups. The most natural generalization is to compute the maximum Chernoff Difference across all groups:
[…]
where the superscript denotes different sensitive groups.
This worst-case Chernoff Difference then captures the largest disparity in classification difficulty across all groups. This corresponds to the maximum notion of equal opportunity:
$\mathrm{EO}_{\max}=\max_{i\neq j\in\mathcal{S}}\lvert\mathrm{FNR}_{i}-\mathrm{FNR}_{j}\rvert$, (29)
which is a common way to define equal opportunity for multi-group settings Hardt et al. (2016); Alghamdi et al. (2022).
@@ Appendix C Isotropic Gaussian Proofs
Lemma 9 (Restatement of Lemma 2).
@@ Appendix D Supplemental Gaussian Figures
D.1 Experimental Details
For the synthetic Gaussian experiments, we use 1-D Gaussian distributions for $P_{0}(x),P_{1}(x),Q_{0}(x),$ and $Q_{1}(x)$ and get 100,000 i.i.d. samples from each distribution to create a balanced dataset. We train a Gaussian Naïve Bayes classifier for each of the groups (this provides the Bayes optimal classifier). We measure the overall accuracy of our classifier and quantify fairness by using the true positive rate disparity between groups (as this corresponds to the dominating exponent). To achieve fairness-accuracy trade-offs, we choose different prior beliefs of our labels for the Gaussian Naïve Bayes classifier. By perturbing the prior probabilities for the unprivileged group, we create fairness-accuracy curves for each of our settings [^6] A proxy for adjusting the threshold of the Bayes optimal classifier (Dutta et al., 2020)..
D.2 Fairness-Accuracy vs Log Fairness-Accuracy
@@
D.3 Case 1: Extended plots
Figure D.4: (Case 1: Extension) $\mu_{0}=0,\mu_{1}=16.5,\sigma=2.43$, $\zeta_{0}=0.5,\zeta_{1}=3.8,\tau=0.55$ (a) Fairness-Accuracy Curve. (b) Log Fairness-Accuracy Curve. After reaching the maximum CD, the fairness-accuracy curves begin to slowly flatten (nearly imperceptible), reflecting the slow decay in CD.
D.4 Case 2: More detail
@@
Figure D.6: (Case 2: Extension) (a) Fairness-Accuracy Curve. (b) Log Fairness-Accuracy Curve. After reaching the reflection point, CD begins to increase. However this spike is so small any change in slope in the fairness-accuracy curve is imperceptible.
@@ E.2 Proofs
Computing the infimum over the open interval (0,1) and closed interval [0,1] yield the same result following the strict convexity of $f$. We point to Remark 2 from Nielsen (2022) for a discussion of the boundary conditions. In fact, in some references, Chernoff Information is presented using this closed interval (Cover and Thomas, 1999). This provides a compactness property that can be utilized in the following lemmas. ∎
Lemma 11.
@@
This follows from Theorem 1 from Newey (1991) as well as pointwise convergence in Lemma 11 and stochastic equicontinuity in Lemma 12. ∎
@@
This result follows from Theorem E.1. Continuity and the uniqueness of the maximizer are given by Nielsen (2022). Compactness is given by Lemma 10. Uniform Convergence in probability is given by Lemma 13. Hence,
@@ F.1.1 Time Complexity
where $n$ is the number of samples used to compute Chernoff Information, $T_{\text{DRE}}$ is time spent on density ratio estimation (DRE) in Line 1, $C_{\text{int}}$ is the time required for one integral operations given in Line 2, $C_{MC}$ is the time required for simple addition/floating point operations required in Line 3, and $T_{\text{cvx}}$ is the time required to solve the convex optimization in Line 4, which is a constant time operation. Hence, for other than $T_{\text{DRE}}$, the rest of the operation is done in $O(n)$ and training a neural network for DRE is often the most time-consuming part, which took approximately 30 minutes on a single NVIDIA L40S GPU.
F.2 Gaussian Estimation
@@ F.3 Datasets
We replicate the previous experiments using 2D embeddings of MNIST data which were obtained via a trained autoencoder with latent dimension $d_{z}=2$.
Table 1: Convolutional autoencoder architecture used for MNIST experiments.
Layer Kernel / Units Stride Output Shape
Input $1\times 28\times 28$
Conv2D + ReLU $3\times 3$, 32 ch. 2 $32\times 14\times 14$
Conv2D + ReLU $3\times 3$, 64 ch. 2 $64\times 7\times 7$
Flatten $64\cdot 7\cdot 7$
Linear $d_{z}=2$ 22
Linear $64\cdot 7\cdot 7$ $64\times 7\times 7$
ConvTranspose2D + ReLU $4\times 4$, 32 ch. 2 $32\times 14\times 14$
ConvTranspose2D + Sigmoid $4\times 4$, 1 ch. 2 $1\times 28\times 28$
We train the autoencoder for 15 epochs, with a batch size of 256, using the Adam optimizer and mean squared error loss function. We choose digit $3$ for $P_{0}$, digit $4$ for $P_{1}$, digit $7$ for $Q_{0}$, and digit $9$ for $Q_{1}$.